The sale of the shoe brand has risen exponentially. One particular shoe store has been experiencing a shape increase in sales and has just hired a graduate, Mr. Perry as Manager. The demand for the brand shoes is 150 boxes per month. Each box cost $ 12,000. The annual holding cost rate is 13%. The particular shoe store ordering cost is $4,300 and the store and warehouse are opened 312 days per year. The lead time for each order is a long 20 days since the shoes are coming all the way from England.


  1. How many boxes of shoes do you recommend that Mr. Perry orders every time he places an order?
  2. Determine the estimated holding cost; and ordering cost.
  3. Would it be true to say it takes Mr. Perry almost 55 days between orders? (Justify your answer.)
  4. How many orders will Mr. Perry have placed each year?
  5. Determine the reorder point.
  6. What can Mr. Perry expect his Total Annual cost to be?
1
Expert's answer
2021-12-07T07:33:03-0500

1.Since the lead time for each order is 20 days, I recommend ordering 300 boxes.

2.Order cost:

"4300\times2=8600"

Storage cost:

"4300\times2+12000\times300=3 608 600"

3.Yes, it looks like a real deadline, since the 20 is the deadline for the order and also the delivery of the goods.

4."\frac{312}{55}=5.67"

6 orders will Mr. Perry have placed each year

5."Reorder Point = Normal consumption during lead-time + Safety Stock=150+150=300"

6."Total Annual cost=12000\\times6\\times150+4300\\times6+0.13(12000\\times6\\times150+4300\\times6)=10\u00a0825\u00a0800+1\u00a0407\u00a0354=12\u00a0233\u00a0154"